Word problem
The population growth of an animal species is described by
F(t) = 700 + 50 log3 (2t + 1),
where t is measured in months. Find the population of this species in an area 13 month(s) after the species is introduced.

Answer by RAY100(1637)

IF,,,f(t) = 700 + 50 *log{3(2t+1)}
.
f(13) = 700 +50*log{3(2(13)+1)},,,,,subst t=13
.
f(13) = 700 + 50*log{3(27)}
.
f(13) = 700 + 50*log {81}
.
f(13) = 700 + 50*1.908
.
f(13) = 700 + 95.424
.
f(13) = 795.424
.